Status Report

ISS On-Orbit Status 27 Jul 2002

By SpaceRef Editor
July 27, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously.  Today the crew completes their 50th day on

The crew is enjoying an off-duty weekend day, with only few routine tasks scheduled (and some items in the "job jar").

As every Saturday, a thorough cleaning of their home kept the station residents busy for three hours.

CDR Valery Korzun completed the daily routine maintenance of the SOSh life
support system.

FE-1 Peggy Whitson received instructions for trying again to clean the heads of the video camcorder.  If the uplinked procedure does not work, she
is to stow the camera.  It will still be available for live events, bu
t can not be used for recording or playbacks until the head problem is resolved.

The crew was also assured that troubleshooting of the TVIS treadmill noise
is a very high priority for ground specialists and that they will work hard
to get IFM (in-flight maintenance) procedures on board.

All crewmembers performed their daily physical exercise, and Valery and Sergei had their periodic PPC (private psychological conference) via VHF radio.

Science Update (Expedition Five — 7th):

The crew was thanked by Lead Increment Scientist Vic Cooley for the SUB
SA troubleshooting and HRF and ADVASC work this week.  The ground team is refining the SUBSA fault tree and running tests on the integrated SUBSA/M
SG ground unit.  It is hoped that on-orbit operations can be resumed next Wednesday with a test run, i.e., without a sample.  The test run would ensure the furnace will rise to the desired temperature.  The ground
will probably re-run the first two samples after the remaining eight samples are processed.  The seed crystal in each of the two processed samples
did not melt, so those samples can be reprocessed without loss.  POC (
Payload Operations Center) gratefully received about two hours of unexpected
live video of the crew’s "operation" of the Education Payload, aka Toys in Space, on Friday.  A copy will be provided to the PI (principal investigator) at the Houston Museum of Natural Science to see if any more demonstrations can be accomplished with the same "toys".

Extra-Vehicular Activity Radiation Monitors (EVARM):  The crew
was thanked for the successful completion of the monthly EVARM background radiation badge reading, data download and downlink.  The data will be provided to the PI for analysis to help characterize on-board radiation level

GASMAP/Pulmonary Function in Flight (PuFF):   The GASMAP 90 Day Health Check was successfully completed this week.  After having
noted that Calibration Module Tank 1 had a >100 psi loss in pressure, it
was determined by the hardware engineers that this loss was acceptable and will not impact future hardware usage.  The crew was thanked again for the great PuFF session this week.  The PuFF files were received on the ground during Thursday’s data downlink and will be forwarded to the PI team.< BR>

Renal (Kidney) Stone Experiment:   Next data collection se
ssion is scheduled for early August.

Interactions (NTXN):  Another big ‘thank you’ from the Interactions team to the crewmembers for their continued participation.

Human Research Facility/Workstation (HRF WS):   n/a

Commercial refrigerator incubator module-commercial Stelsys (STELSYS): &
Experiment is complete and samples are stowed in frozen state in AR
CTIC-1. "Waiting for a ride home".

Microencapsulation Electrostatic Processing (MEPS):
Thanks to the crew from the MEPS team, which can92t wait for PCM (processing chamber module) return on 9A.

Advanced Astroculture (ADVASC):  The ADVASC team thanked Peggy
Whitson for all her great work with ADVASC this week in completing the 2 Plant Photos activities and the Condensate and Nutrient Samples.  After viewing yesterday’s photos the ADVASC team has confirmed that the soybeans have flowered, and the team will soon be adjusting the environmental conditions
to foster the seed development. The ADVASC hardware has continued to operate nominally.

Solidification Using a Baffle in Sealed Ampoules (SUBSA):  The
SUBSA runs scheduled this past week did not go as planned. It appears there
is a software bug which has prevented commanding to operate the furnace to the temperatures required to achieve the science. The furnace controller software has been updated and is being tested with the SUBSA hardware in the MSG
engineering unit on the ground. When that testing has been successfully completed and verified, the new software will be uplinked. If the ground can work them in to the timeline, SUBSA samples 6 and 10 will be rerun, and it is
expected that the originally planned science can be accomplished.

Pore Formation and Mobility Investigation (PFMI):  New payload
for Increment 5 will be performed during Stage 9A.

Space Acceleration Measurement System (SAMS):  SAMS/PIMS: SAMS
heads are nominal and collecting acceleration data for general characterization of the microgravity environment in the vibratory regime.  SAMS head
121f08 was activated in support of SUBSA run and initial characterization of the environment in the MSG rack. Currently ground processing is working on
identifying many intriguing signatures captured in the previous weeks, including those from GASMAP and MEPS-PCM operations.

Microgravity Acceleration Measurement System (MAMS):  MAMS is active and the OARE Sensor Subsystem (OSS) continues to collect quasi-steady
data for general characterization of this important low-frequency, relatively low-amplitude acceleration regime and in support of science payloads currently conducting operations.  HiRAP data transmission is disabled.

Active Rack Isolation System (ARIS) Support:  This week, ARIS successfully completed characterization of the foam snubber isolators.  
These foam inserts, along with ARIS snubber assemblies already on orbit, are
being considered for installation on the HRF rack to help it meet its vibration requirements.  With the ARIS alignment guides installed on the ARIS rack, a couple of isolation tests will be run to compare the

vibration environment on the ARIS rack when it is locked down with the guides as compared to when ARIS is commanded to a position HOLD control mode. &n
bsp;Peggy was thanked for her answers to  questions on the installation
of the alignment guides.  They gave valuable insight as the ground considers redesign options to make the alignment guides easier to use.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES)
 The STES-8 unit is running nominal at 22.0 deg C since transfer.

Materials ISS Experiment (MISSE):  In progress. Deployed outsid
e. Nominal and collecting data.

Educational Payload Operations (EPO):  EPO-5 thanked the crew for completing their International Toys in Space Payload objectives.  The addition of the Ku-band was great and the footage looked fantastic.  It looked like the crew was having as much fun doing the demonstrations as the ground was having watching the Ku-band feed.  No issues or concerns.
 Preliminary indications show that all objectives were met.

EarthKAM:   Planned.  Will be performed during Stage

All EarthKAM images are available for public access on the EarthKAM data system at:

Zeolite Crystal Growth (ZCG):  Plans are being considered to possibly put seven of the metal autoclaves in ARCTIC-2.  ZCG would like to use ARCTIC 2 to attempt to slow down the degradation of the science samples due to the extended stage.

ARCTIC Refrigerator/Freezer 1 (ARCTIC-1):  the ARCTIC-1 front panel LCD (liquid crystal display) is generating "garbage" characters instead of numbers, but this is no impact to science or safety, and troubleshooting of this issue is underway.

ARCTIC Refrigerator/Freezer 2 (ARCTIC-2):  As a precautionary measure, the current thinking is that the planned first round of troubleshooting to improve water flow to ARCTIC-2 will not take place until StelSys has
been removed from ARCTIC-1.

Commercial Generic Bioprocessing Apparatus (CGBA):  Will be per
formed during Stage 9A.

Crew Earth Observations (CEO): The ground received 213 new images downlinked from ISS this week, bringing the total to 2,630 to date.  This batch included a number of extraordinary views of fires and burns scars in
our target area of Zambia.  Thanks to the crew for their continued efforts during the poor lighting of the Southern Hemisphere winter and the attitude limitations of XPOP.

SpaceRef staff editor.